2. On February 1, 2018, Henry Corporation purchased 200,000 shares of Glover, Inc., common stock for $5,400,000. The investment represents a 40% interest in the net assets of Glover and gave Henry the ability to exercise significant influence over Glovers operations On September 30, 2018, Henry received dividends of $1.00 per share on the Glover investment. Glover reported net income of $300,000 for the year ended December 31, 2018. The market value of the Glover stock was $30 per share on December 31, 2018. 9199 On February 1, 2018, the book value of Glovers net assets was $8,000,000. The fair value of Glovers depreciable assets, with an average remaining useful life of ten years, exceeded their book value by $800,000. Glover depreciates its assets by the straight-line method. The remainder of the excess of the cost of the investment over the book value of net assets purchased was attributable to goodwill. INSTRUCTIONS a) Prepare all appropriate journal entries related to the investment during 2018, assuming Henry accounts for this investment by the equity method. Prepare all appropriate journal entries related to the investment during 2018, assuming that the 200,000 shares of Glover represent a 5% interest in the net assets of Glover rather than a 40% interest. b)
Answer #1
a) Journal Entries:
Date Accounts Title Debit $ Credit$
2018
1-Feb Equity Investment in Glover 5400000
Cash 5400000
(Being equity investment with 40% in Glover co's
200000 common stock)
30-Sep Cash 200000
Equity Investment in Glover 200000
(Being dividend recd. @ $1 on 200000 common stock
of Glover Inc.)
31-Dec Equity Investment in Glover
Net Income of Glover
(Being 40% share of Glover's 2018 earned booked)
31-Dec Depreciation-Goodwill 220000
Equity Investment in Glover 220000
(Being Excess of book value is treated as Goodwill
and depreciated on SLM with 10 years life)
(5400000-8000000*0.40)/10=$220000
b) Journal Entries:
Date Accounts Title Debit $ Credit$
2018
1-Feb Available for Sale- Glover 5400000
Cash 5400000
(Being investment with5% in Glover co's
200000 common stock)
30-Sep Cash 200000
Dividend Income 200000
(Being dividend recd. @ $1 on 200000 common stock
of Glover Inc.)
31-Dec Available for Sale- Glover 600000
Unrealised Income on re-valuation 600000
(Being revaluation of investment at year end when
Glover share market value was $30 per share
200000*30 - 5400000 = 600000)
(No Goodwill depreciation & No annual income transfer entry required
if total investment is 5% of total capital of Glover)